The seismicity during today has been at about the same level as yesterday, and dominated by long-period signals. Fifty-seven such signals were recorded, of which about 25 were due to rockfalls from the dome. Two small volcano-tectonic earthquakes were also recorded, one at a depth of about 3 km beneath Hermitage and the other at a depth of 2 km, east of St George's Hill. A signal was also recorded from a regional earthquake, which was more than 500 km away.
The eastern EDM triangle was measured today, and showed almost no movement of Castle Peak since it was last measured yesterday.
Visual observations were made this morning from the helicopter and from Whites and Long Ground, in generally cloudy conditions. The dome continues to steam vigorously, and the most active area appears to be in the east. Over the last two days, rockfalls have occurred from two parts of the dome, near Farrell's and to the north of Castle Peak. In the latter area, material has fallen down a narrow channel towards the Tar River Soufriere. These rockfalls have been the source of the frequent ash clouds observed recently. Observers at Whites this morning observed almost continuous ash and steam production.
Dr William Ambeh arrived back in Montserrat last night. He will be taking over as Chief Scientist from Tuesday.